OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation 3x − 2y = 6 a= x-intercept: (0, −3), y-intercept: (0, 2) b= x-intercept: (−3, 0), y-intercept: (2, 0) c= x-intercept: (0, 2), y-intercept: (0, −3) d= x-intercept: (2, 0), y-intercept: (0, −3)

OpenStudy (saifoo.khan):

When x intercpts y = 0, when y intercpts x = 0

OpenStudy (anonymous):

hmm....?

OpenStudy (saifoo.khan):

Alright, so: x intercpt, \[3x − 2y = 6\]\[3x − 2(0) = 6\]\[x =2\]

OpenStudy (saifoo.khan):

When y intercpts, x = 0 \[3x − 2y = 6\]\[3(0) − 2y = 6\]\[y = -3\]

OpenStudy (saifoo.khan):

Therefore: x = (2, 0), y = (0,-3)